Grade 11Mechanics

A particle revolves with constant angular acceleration pi rad / sec.. If the particle starts from rest, how many revolutions will it make in 10 sec is

Ravichodipillli
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear Student,
This means
Particle rotates pi radian in 1 sec
Hence it will rotate 10\pi radian in 10 seconds
Now
1 revolution = 2\pi
Thus
Total number of revolutions will be = 10\pi/2\pi = 5
Hence the answer will be 5.
 
Regards,
Arun(AskIITians Forum Expert)